### v1

# Plan-R1: Safe and Feasible Trajectory Planning as Language Modeling

Abstract
Safe and feasible trajectory planning is essential for real-world autonomous driving systems. However, existing learning-based planning methods often rely on expert demonstrations, which not only lack explicit safety awareness but also risk inheriting unsafe behaviors such as speeding from suboptimal human driving data. Inspired by the success of large language models, we propose Plan-R1, a novel two-stage trajectory planning framework that formulates trajectory planning as a sequential prediction task, guided by explicit planning principles such as safety, comfort, and traffic rule compliance. In the first stage, we train an autoregressive trajectory predictor via next motion token prediction on expert data. In the second stage, we design rule-based rewards (e.g., collision avoidance, speed limits) and fine-tune the model using Group Relative Policy Optimization (GRPO), a reinforcement learning strategy, to align its predictions with these planning principles. Experiments on the nuPlan benchmark demonstrate that our Plan-R1 significantly improves planning safety and feasibility, achieving state-of-the-art performance.
1 Introduction
Trajectory planning is a fundamental component of autonomous driving systems, directly influencing vehicle safety, efficiency, and the ability to navigate complex and dynamic environments. In recent years, learning-based planning approaches have attracted increasing attention due to their strong adaptability, competitive performance, and minimal reliance on manually designed rules. These methods offer promising solutions for generating high-quality context-sensitive trajectories that can respond effectively to various traffic scenarios and rapidly changing road conditions.
However, due to the complexity of large-scale real-world datasets such as nuPlan [2], most existing planning methods, whether based on imitation learning [36, 5, 4, 11, 39, 17, 8] or reinforcement learning [37, 9, 12, 15], rely heavily on expert demonstrations for supervision. This dependency introduces two key limitations. First, expert data often lacks negative scenarios (e.g., collisions, off-road driving), making it difficult for models to learn explicit safety constraints such as maintaining drivable area compliance or avoiding collisions. Second, expert demonstrations are not always optimal and may contain unsafe behaviors. For example, we find that approximately 11.2% of the nuPlan training scenes exhibit speeding, and some contain uncomfortable maneuvers or critically low time-to-collision (TTC) values. Models trained directly on such data risk inheriting these undesirable behaviors without a clear notion of safety or feasibility.
To address these limitations, we draw inspiration from the success of large language models (LLMs) [29, 21], which typically follow a two-stage training paradigm. LLMs are first trained as a general-purpose predictor via self-supervised pretraining, and then fine-tuned using reinforcement learning from human feedback (RLHF) to align outputs with human preferences. Our key insight is that trajectory planning can similarly be formulated as a two-stage training process, i.e., first trained as a sequence prediction model on expert driving data, and then fine-tuned using reinforcement learning to align the planning behavior with explicit planning principles, such as safety, comfort, and rule compliance. This decoupled paradigm separates behavior learning and planning principle alignment, enabling the system to retain valuable driving behaviors while discarding unsafe patterns learned from the expert data.
Specifically, our proposed Plan-R1 method consists of two sequential stages. In the pre-training stage, the trajectories are discretized into motion tokens across time and space following prior studies [28, 35, 23], and an autoregressive model is trained via next motion token prediction. In the fine-tuning stage, the pre-trained model is frozen as a reactive world model to predict the motions of surrounding agents, while a duplicate copy is fine-tuned for ego trajectory planning. These two models collaboratively perform multi-step predictions to generate the joint future of the ego vehicle and surrounding agents. To guide fine-tuning, we design interpretable, rule-based reward functions that evaluate various aspects of planned trajectories, including but not limited to on-road adherence, collision avoidance, speed limit compliance, and driving comfort. The ego planner is then fine-tuned using Generalized Reinforcement Policy Optimization (GRPO) [29] algorithm with these rewards. Experiments on the nuPlan benchmark demonstrate that Plan-R1 significantly enhances planning safety and feasibility, achieving state-of-the-art (SOTA) performance.
The primary contributions of this paper are:
  • We propose a new perspective that formulates trajectory planning as a principle-aligned sequence prediction task, enabling the decoupling of behavior learning and planning principle alignment.
  • We introduce Plan-R1, a two-stage trajectory planning framework that combines autoregressive pre-training with reinforcement learning fine-tuning to align with planning principles.
  • Instead of relying on preference data for fine-tuning, a set of interpretable, rule-based reward functions is designed to capture essential planning principles, including but not limited to drivable area compliance, collision avoidance, speed limit compliance, and driving comfort.
  • Our Plan-R1 significantly improves the safety and feasibility of planned trajectories, achieving SOTA performance on the nuPlan benchmark, particularly outperforming existing planners in the reactive closed-loop simulation setting.
2 Related Work
2.1 Learning-based trajectory planning
Learning-based trajectory planning methods for autonomous driving can be broadly categorized into imitation learning (IL) and reinforcement learning (RL). IL trains policies to imitate expert behavior, typically using large-scale datasets of human driving trajectories, and has been shown to produce realistic, human-like motion in complex urban environments [1, 26, 5, 4, 24, 34]. RL, on the other hand, optimizes policies through interaction with the environment to maximize task-specific rewards, enabling the discovery of behaviors beyond those observed in demonstrations [37, 15, 16, 12, 9].
Despite promising results in simulation, both IL and RL face key limitations when applied to real-world planning. IL is fundamentally constrained by the quality and coverage of expert demonstrations. RL methods, while more flexible, often address the challenges of complex reward design and sample inefficiency by using the negative distance to expert trajectories as the reward function [37, 15], thereby remaining dependent on expert data. However, expert trajectories are not always safe or optimal, and models trained on such data may inherit undesirable behaviors, such as traffic violations or infeasible maneuvers. Our key insight is that explicitly decoupling behavior learning from planning principle alignment can help mitigate this issue—a direction largely overlooked in prior IL- and RL-based planners.
2.2 Trajectory generation as language modeling
Inspired by the success of LLMs, recent studies have explored formulating trajectory generation as a sequence modeling task. These approaches [28, 35, 23] discretize agent trajectories into motion tokens and train autoregressive models via next motion token prediction. This formulation allows the model to learn the multi-modal distribution of driving behaviors from demonstrations and has shown strong performance in trajectory prediction and simulation. However, most of these methods focus exclusively on the pre-training stage, overlooking the fine-tuning phase that is now standard in modern LLM pipelines, such as RLHF in InstructGPT [21]. While these models can produce diverse and plausible trajectories, they often lack guarantees of safety, feasibility, or rule compliance during execution. Our work bridges this gap by introducing reinforcement learning-based fine-tuning to align generated trajectories with explicit planning principles, analogous to how LLMs are aligned with human preference, but tailored to safety-critical planning.
Concurrent work TrajHF [15] and earlier Gen-Drive [12] both align planning behavior with human preferences via reinforcement learning from human feedback (RLHF). In contrast to our use of interpretable rule-based rewards derived from domain knowledge, these methods depend heavily on preference data. TrajHF minimizes the distance to expert-annotated preferred trajectories, while Gen-Drive trains a reward model from pairwise comparisons to guide learning. Both face two key limitations: collecting high-quality preferences is costly and labor-intensive, and performance is highly sensitive to data quality and coverage, limiting generalization beyond the annotated distribution.
2.3 Reinforcement learning algorithms for sequence-level fine-tuning
Recent progress in LLMs has demonstrated the effectiveness of reinforcement learning algorithms in aligning sequence generation with desired objectives. In particular, Proximal Policy Optimization (PPO) [27] has emerged as the practical choice for RLHF, and has been widely adopted for fine-tuning LLMs such as InstructGPT [21]. More recently, GRPO [29] has been proposed as a lightweight alternative that eliminates the need for value functions by computing relative advantages within a group of sampled sequences. Compared to PPO, GRPO simplifies implementation by removing the value critic, reducing memory overhead, and avoiding instability caused by inaccurate value estimates. GRPO has demonstrated strong performance across various LLM tasks involving search [18, 10] and reasoning [22, 13]. Motivated by the sequential nature of trajectory generation, which closely parallels text generation, we adopt GRPO to fine-tune our planning model. Our experiments demonstrate that this formulation enables reinforcement learning to effectively align motion sequence generation with high-level planning principles.
3 Method
Figure 1: Illustration of our two-stage Plan-R1. In the pre-training stage, an autoregressive model is trained via next motion token prediction on expert data. In the fine-tuning stage, the model is optimized using the GRPO algorithm with rule-based rewards to align the generated trajectories with planning principles such as safety, feasibility, and comfort.
We propose Plan-R1, a two-stage framework that decouples behavior learning and planning principle alignment to improve trajectory safety and feasibility. As illustrated in Figure 1, Plan-R1 consists of an autoregressive pre-training stage and a reinforcement learning fine-tuning stage. The model is first trained on expert demonstrations to learn multimodal motion behaviors, and then fine-tuned with rule-based rewards to align the ego vehicle’s behavior with explicit planning principles. In Section 3.1, we present our formulation of trajectory planning as a principle-aligned sequence generation task. In Section 3.2, we describe the autoregressive pre-training process on expert data. Finally, in Section 3.3, we detail how fine-tuning is performed via the GRPO algorithm with trajectory-level rewards.
3.1 Problem formulation
Trajectory generation is inherently sequential, which makes autoregressive modeling a natural and effective choice for this task. Building on this formulation, we further incorporate alignment with high-level planning principles such as safety, comfort, and traffic rule compliance. Concretely, we reformulate trajectory planning as a principle-aligned autoregressive generation task. In this formulation, the ego vehicle’s trajectory is generated jointly with the predicted motions of surrounding agents [7, 38, 3], enabling the planner to account for multi-agent interactions and ensure safe, feasible planning. Specifically, given a scene context $C$—including current and historical observations, map information, and navigation routes—and a set of planning principles $P$, the goal is to generate the joint future trajectories for all agents:
$Y=f(C,P),\quad Y=\{y_{t,n}\}_{\ t\in[1:F],n\in[0:N]},$ (1)
where agent 00 is ego, $N$ denotes the number of surrounding agents and $F$ is the planning horizon. Directly modeling the joint distribution $p(Y\mid C,P)$ is intractable due to its high dimensionality. However, by leveraging temporal causality [28] and the short-term conditional independence between agents [31, 25], the joint distribution can be factorized as:
$p(Y\mid C,P)$ $=\prod_{t=1}^{F}p(y_{t,0:N}\mid y_{<t,0:N},C,P)$ (2)
$=\prod_{t=1}^{F}p(y_{t,0}\mid y_{<t,0:N},C,P)\prod_{n=1}^{N}p(y_{t,n}\mid y_{<t,0:N},C,P)$
$\approx\prod_{t=1}^{F}\pi_{e}(y_{t,0}\mid y_{<t,0:N},C,P)\prod_{n=1}^{N}p_{a}(y_{t,n}\mid y_{<t,0:N},C),$
where the final line reflects that the behaviors of surrounding agents are independent of the ego’s planning principles. This decomposition yields two sub-problems: (i) $p_{a}(y_{t,n}\mid y_{<t,0:N},C)$, predicting each surrounding agent’s motion conditioned on the joint history and scene context, and (ii) $\pi_{e}(y_{t,0}\mid y_{<t,0:N},C,P)$, generating the ego trajectory with additional conditioning on planning principles. The distribution $p_{a}$ is naturally modeled as an autoregressive sequence prediction task, in which the future motion of each agent is inferred from its own history and the history of surrounding agents, along with the scene context. $\pi_{e}$ introduces additional constraints on the ego trajectory to meet planning principles such as safety, comfort, and rule compliance. Since $p_{a}$ is trained on human driving data, which inherently follows these planning principles, it captures a range of plausible and realistic behaviors. Therefore, we assume that $p_{a}$ provides a reasonable prior for $\pi_{e}$. However, $p_{a}$ lacks explicit alignment with these planning objectives, and its generated trajectories may still violate safety constraints or result in suboptimal decisions in complex scenarios, rendering it unsuitable for direct deployment in safety-critical planning tasks. To address this, we initialize $\pi_{e}$ with $p_{a}$, then fine-tune $\pi_{e}$ using reinforcement learning with rule-based rewards to explicitly align the ego’s behavior with the planning principles.
3.2 Autoregressive pre-training
Tokenization. To facilitate sequence modeling, continuous trajectories are discretized into motion tokens. Temporally, trajectories are segmented at fixed intervals. Spatially, the K-disk clustering algorithm [23] is applied to the resulting motion segments based on their average corner distance, producing a motion token vocabulary. Each token represents a prototypical displacement and heading change over a fixed time step. For each agent, the entire trajectory is thus transformed into a sequence of such motion tokens spanning the prediction horizon.
Model architecture. Our model employs a transformer decoder with factorized attention [19, 20] to encode multi-agent spatiotemporal interactions, as shown in Figure 1 (a). It is designed to capture the temporal dynamics of each agent’s trajectory and its interactions with the map and nearby agents.
Following prior work [35, 28], the model processes motion tokens through a stack of attention-based fusion blocks at each time step. Each block sequentially applies three key components: temporal self-attention, agent-map cross-attention, and agent-agent cross-attention. Temporal self-attention models the motion continuity of an individual agent over time, enabling the model to capture temporal dependencies in the agent’s trajectory. Agent-map cross-attention incorporates spatial constraints by attending to relevant road elements, guiding trajectories to comply with the road network. Finally, agent-agent cross-attention captures local interactions among neighboring agents within a fixed distance threshold, enabling the model to reason about agent interactions in dynamic environments.
To preserve translation invariance and rotation invariance, relative spatiotemporal position embedding [41, 32] is applied to all attention modules. The ego and non-ego agents share the same encoder structure, and all attention computations are performed in a query-centric manner to support efficient multi-agent rollout.
Training objective. The model is trained with a next motion token prediction objective. Given a tokenized trajectory sequence, the model is trained to minimize the negative log-likelihood of the next motion token at each step:
$\mathcal{L}_{pretrain}=-\sum_{t=1}^{F}\sum_{n=0}^{N}\log p_{a}(a_{t,n}|a_{<t,0:N},C),$ (3)
where $a_{t,n}$ is the motion token of agent $i$ at time $t$. This objective allows the model to approximate the latent distribution of human driving behaviors, capturing diverse and realistic motion patterns as demonstrated in large-scale expert data. Notably, no explicit planning principles are introduced at this stage; the model purely imitates human-like behavior conditioned on the scene context.
3.3 Reinforcement learning fine-tuning
While autoregressive pre-training captures human-like motion patterns, it does not explicitly enforce adherence to planning principles. This is particularly problematic when expert trajectories include undesirable behaviors—e.g., speeding or uncomfortable maneuvers—which may be inadvertently retained during pretraining. To mitigate this issue, we fine-tune the ego trajectory generator using reinforcement learning, as shown in Figure 1 (b), optimizing it to align with predefined planning principles. The generation process is framed as a sequential decision-making problem, where the ego agent selects motion tokens to maximize rewards that reflect desirable planning behaviors.
Dual-Model rollout. To enable stable multi-agent rollout and principle-aligned fine-tuning, we adopt a dual-model design. Specifically, we freeze the pretrained autoregressive model for predicting the future motions of the surrounding agents, and create a trainable replica to represent the ego policy $\pi_{e}$. During each rollout, the two models jointly generate trajectories step by step. The frozen model, serving as a reactive world model, predicts the future motions of surrounding agents based on the history of all agents, including the ego. Although fixed during training, the world model remains responsive to ego actions by updating its predictions based on the ego’s historical motion tokens. This interaction ensures that the surrounding agents’ behaviors are aligned with the ego vehicle’s actions, ensuring realistic, interaction-aware simulation in multi-agent settings. Meanwhile, the trainable ego model explores alternative actions and receives feedback through the reward function. This separation enables ego-centric policy updates without destabilizing surrounding agent dynamics, while also ensuring that all agents exhibit natural reactive behaviors, as observed in real-world interactions.
Rule-based rewards. To align the ego policy with planning principles, we design a set of interpretable, rule-based reward functions. Rewards are computed based on the final predicted trajectories and are designed to reflect key planning properties such as safety, feasibility, and comfort. Following [2, 9], we compute the total reward as the product of two components: a set of multiplicative safety indicators (e.g., driving area compliance) and a weighted sum of soft cost terms (e.g., comfort):
$R(Y)=\prod_{k\in\mathcal{I}_{\text{safe}}}\mathbf{1}_{k}\cdot\sum_{j\in\mathcal{I}_{\text{cost}}}w_{j}\cdot r_{j}(Y),$ (4)
where $\mathbf{1}_{k}\in\{0,1\}$ denotes whether safety constraint $k$ is satisfied, and $r_{j}(Y)$ is the score of cost term $j$ with weight $w_{j}$. This formulation ensures that violations of critical safety conditions will nullify the total reward, while soft planning objectives are optimized when safety constraints are satisfied. Details of the reward components and implementation are provided in the Appendix.
Optimization with GRPO. To fine-tune the ego policy toward planning principles, we adopt GRPO as our reinforcement learning algorithm. Compared to PPO, GRPO eliminates the need for an explicit value function by computing relative advantages within sampled trajectory groups. This reduces implementation complexity, avoids the instability associated with inaccurate value estimates, and improves sampling efficiency, making it well-suited for trajectory-level planning.
During fine-tuning, we treat the pre-trained trajectory predictor as the reference policy $\pi_{\text{ref}}$, and the ego trajectory generator as the trainable policy $\pi_{e}$. For each scenario, GRPO samples a group of joint future trajectories $\{Y_{1},Y2,\dots,Y_{G}\}$ from the old policy $\pi_{e_{\text{old}}}$, where each sample contains both ego and nearby agents’ predicted trajectories. The trainable policy $\pi_{e}$ is updated by minimizing this loss:
$\mathcal{L}_{\text{finetune}}=-\frac{1}{G}\sum_{g=1}^{G}\left(\frac{\pi_{e}(Y_{g}\mid C,P)}{\pi_{e_{\text{old}}}(Y_{g}\mid C,P)}\hat{A}_{g}-\beta\ D_{\text{KL}}\left[\pi_{e}\,\|\,\pi_{\text{ref}}\right]\right),$ (5)
where $\hat{A}_{g}$ is the normalized relative advantage of the $g$-th joint future $Y_{g}$ within the sampled group. $\hat{A}_{g}=(R(Y_{g})-\mu_{R})/\sigma_{R}$, where $\mu_{R}$, $\sigma_{R}$ are the mean and standard deviation of the group’s reward values, respectively. This group-wise normalization eliminates the need for a learned value critic, allowing stable optimization through relative comparison across trajectories in the same rollout batch.
The KL divergence term in the loss serves as a regularizer, constraining the updated policy $\pi_{e}$ to remain close to the reference policy $\pi_{\text{ref}}$. This helps preserve the prior learned during pretraining, preventing the fine-tuned policy from deviating too far from reasonable behaviors.
4 Experiments
4.1 Experimental setup
Datasets. We evaluate our method on the nuPlan [2] benchmark, a large-scale platform designed for trajectory planning in autonomous driving. The dataset contains over 1,300 hours of expert driving logs collected across four cities, covering diverse real-world scenarios, such as merges, intersections, roundabouts, lane changes, and interactions with pedestrians. The benchmark supports simulation-based evaluation, where scenarios are replayed for 15 seconds at 10 Hz. The ego vehicle executes planned trajectories via a bicycle model and an LQR controller, while other traffic agents follow either replayed or Intelligent Driver Model (IDM) [33] policies. Following PLUTO [4], we sample 1M training instances across all scenario types for autoregressive pretraining. To reduce the computational cost of closed-loop rollouts during reinforcement learning, we construct a smaller dataset of 100,000 scenarios for fine-tuning. Evaluations are conducted on the standard Val14 [6], Test14-random, and Test14-hard [5] splits under both non-reactive and reactive simulation settings.
Metrics. We evaluate planning performance in closed-loop simulation using the Non-Reactive Closed-Loop Score (NR-CLS) and Reactive Closed-Loop Score (R-CLS). These metrics assess the planner’s ability to generate safe and feasible trajectories over 15-second rollouts, considering collisions, off-road driving, speed limit violations, and the progress of the expert route. In the NR-CLS setting, surrounding agents follow replayed logged trajectories, while in R-CLS, they react using the IDM [33] planner. Each score ranges from 0 to 100, where a higher score indicates better performance.
Table 1: Comparison with SOTAs on nuplan benchmark. The best result is in bold and the second best result is underlined. *: with rule-based post-processing. NR/R: non-reactive/reactive mode.
Type Planner Val14 Test14-hard Test14-random
NR R NR R NR R
Expert Log-Replay 93.53 80.32 85.96 68.80 94.03 75.86
Rule-based & Hybrid IDM [33] 75.60 77.33 56.15 62.26 70.39 72.42
PDM-Closed* [6] 92.84 92.12 65.08 75.19 90.05 91.64
PDM-Hybrid* [6] 92.77 92.11 65.99 76.07 90.10 91.28
Gameformer* [11] 79.94 79.78 68.70 67.05 83.88 82.05
PLUTO* [4] 92.88 89.84 80.08 76.88 92.23 90.29
PlanAgent* [40] 93.26 92.75 72.51 76.82 - -
Diffusion Planner* [39] 94.26 92.90 78.87 82.00 94.80 91.75
Carplanner* [37] - - - - 94.07 91.10
Plan-R1* (Ours) 94.77 93.25 77.79 80.97 94.11 94.51
Learning-based UrbanDriver [26] 68.57 64.11 50.40 49.95 51.83 67.15
PDM-Open [6] 53.53 54.24 33.51 35.83 52.81 57.23
PlanTF [5] 84.27 76.95 69.70 61.61 85.62 79.58
PLUTO [4] 88.89 78.11 70.03 59.74 89.90 78.62
Diffusion Planner [39] 89.87 82.80 75.99 69.22 89.19 82.93
Plan-R1 (Ours) 87.99 84.97 75.31 73.18 89.46 88.89
(a) Expert
(b) Pluto
(c) Diffusion Planner
(d) Plan-R1 (Ours)
Figure 2: Comparison of closed-loop ego trajectories in a non-reactive scenario without post-processing. Trajectories are color-coded: orange marks speeding segments, while green indicates compliant motion. The expert trajectory (a) contains a clear speeding segment, replicated by both PLUTO (b) and Diffusion Planner (c), showing that the planners based on expert data inherit undesirable behaviors therein. In contrast, Plan-R1 (d) avoids speeding, demonstrating the effectiveness of reinforcement learning fine-tuning in promoting safety-compliant behavior.
4.2 Comparison with SOTAs
Quantitative results. We compare our method, Plan-R1, against a wide range of existing planners on the nuPlan benchmark, as shown in Table 1. Across all three benchmarks, Plan-R1 achieves performance comparable to the strongest prior method, Diffusion Planner [39], under the non-reactive setting. This indicates that our reinforcement learning fine-tuning preserves the model’s ability to align with expert-like behavior. Under the more challenging reactive setting, where surrounding agents respond dynamically to the ego vehicle, Plan-R1 substantially outperforms all baselines. It achieves state-of-the-art reactive scores of 84.97 on Val14, 73.18 on Test14-hard, and 94.51 on Test14-random, surpassing the next-best method, Diffusion Planner, by +2.17, +3.96, and +5.96 points, respectively. Compared to other learning-based planners such as PlanTF [5], Plan-R1 demonstrates significantly stronger generalization to interactive scenarios. We attribute this improvement to the exploration capability introduced by reinforcement learning, which enables the planner to generate safer and more feasible trajectories beyond what is directly covered in the expert demonstrations. Following Diffusion Planner [39], we apply the existing refinement module [30] as a post-processing step to Plan-R1, ensuring a fair comparison with prior methods that adopt similar pipelines. With post-processing, Plan-R1 continues to achieve state-of-the-art performance, reaching a reactive score of 94.51 on Test14-random, outperforming the second-best method by 2.76 points. This result highlights the effectiveness of our approach in producing high-quality trajectories.
We attribute these gains to our two-stage training framework: autoregressive pretraining establishes a strong behavioral prior from expert data, while reinforcement learning fine-tuning explicitly aligns trajectory generation with safety and feasibility objectives. Together, these stages enable Plan-R1 to produce safe, feasible, and socially compliant behaviors in closed-loop planning scenarios.
Qualitative Results. To further illustrate the practical benefits of our approach, we conduct a case study comparing closed-loop non-reactive rollouts across different planners on a representative scenario from the nuPlan dataset. In this scenario, the expert trajectory exhibits a speeding behavior that violates the local speed limit. We evaluate three learning-based planners—PLUTO [4], Diffusion Planner [39], and Plan-R1—none of which incorporate rule-based post-processing in this analysis. As shown in Figure 2, both PLUTO and Diffusion Planner, which are trained solely via supervised learning on expert demonstrations, replicate the speeding behavior in their closed-loop rollouts. In contrast, Plan-R1 successfully avoids speeding and maintains a compliant velocity profile throughout the rollout. This demonstrates our model’s ability to identify and correct undesirable behaviors inherited from expert data, a capability enabled by reinforcement learning fine-tuning with rule-based reward supervision. These results highlight the importance of explicit objective alignment in ensuring rule-compliant and safe planning behavior.
Table 2: The importance of the fine-tuning stage.
Planner Score Collisions TTC Drivable Speed Comfort Progress
Baseline (w/o finetning) 82.81 94.44 90.04 94.64 97.83 99.62 85.71
PlanR1 (w/ finetning) 88.89 96.55 94.25 97.70 99.46 99.46 88.69
$\Delta$ +6.08 +2.11 +4.21 +3.06 +1.63 -0.16 +2.98
Off road Refer to caption
speeding Refer to caption
Collision Refer to caption
(a) Baseline (w/o fine-tuning)
(b) Plan-R1 (w/ fine-tuning)
Figure 3: Comparison of closed-loop ego trajectories generated by the baseline (a) and Plan-R1 (b) in three non-reactive closed-loop simulation scenarios. The baseline exhibits issues such as off-road driving (left), speeding (middle), and collision with static obstacles (right), while Plan-R1 avoids these issues, generating safe and feasible trajectories. orange indicates segments where the ego exceeds the local speed limit, while cyan indicates static obstacles.
4.3 Ablation studies
The importance of the fine-tuning stage. To evaluate the effectiveness of the fine-tuning stage in our Plan-R1 framework, we compare Plan-R1 with a baseline model that undergoes only pretraining, as shown in Table 2. Compared to the baseline without fine-tuning, Plan-R1 shows substantial improvements across key metrics in the reactive closed-loop simulation on Test14-random. Specifically, it achieves a score increase of +6.08, with notable gains in collision avoidance (+2.11), time-to-collision (TTC) (+4.21), drivable area compliance (+3.06), and speed limit compliance (+1.63). Although the comfort score fluctuated slightly (-0.16), the progress score improved significantly (+2.98), suggesting that the model enhances safety and feasibility not by relying on remaining stationary or abrupt braking, but by generating better trajectories.
To further illustrate the benefits of fine-tuning, we compare the closed-loop ego trajectories from the baseline and Plan-R1 in three non-reactive scenarios. As shown in Figure 3, the baseline exhibits off-road driving, speeding, and collision with an obstacle, respectively. In contrast, Plan-R1 avoids these issues, producing trajectories that adhere to the road, obey speed limits, and avoid obstacles. These results highlight the crucial role of fine-tuning in improving planning safety and feasibility.
Figure 4: Effect of KL regularization weight on NR-CLS. A moderate KL weight ($\beta{=}0.1$) achieves the best performance.
Figure 5: Comparison of pass@$k$ between the pretrained baseline and Plan-R1. Plan-R1 always achieves higher pass@$k$.
Effect of KL regularization. We investigate the impact of the KL divergence regularization weight $\beta$ in Equation 5 during the fine-tuning stage. As shown in Figure 5, when $\beta$ is set too small (e.g., 0.01), the updated policy deviates excessively from the pretrained reference, leading to a degradation in performance due to the loss of learned motion priors. Conversely, when $\beta$ is too large (e.g., 1.0), the policy is overly constrained to remain close to the reference model, limiting the benefits of fine-tuning and resulting in similar results. A moderate value (e.g., $\beta{=}0.1$) achieves the best balance, preserving the distributional prior while enabling meaningful adaptation toward planning objectives.
Pass@$k$ analysis. To further assess the reliability of our method under varying sampling conditions, we conduct a pass@$k$ [14] analysis on the validation set. For both the pretrained baseline and our Plan-R1, we generate $k$ trajectories per scenario and compute the proportion of cases where at least one satisfies all core safety and feasibility criteria—i.e., staying within drivable areas, avoiding collisions, respecting speed limits, and maintaining comfort. As shown in Figure 5, pass@$k$ improves with increasing $k$ for both models. However, Plan-R1 consistently outperforms the baseline, indicating it yields high-quality trajectories with greater likelihood. This suggests that fine-tuning not only improves the single-shot performance but also enhances the model’s ability to generate feasible trajectories across multiple samples, which is critical for robust planning under uncertainty.
5 Conclusion
In this work, we introduced Plan-R1, a novel two-stage framework for safe and feasible trajectory planning. Inspired by the success of LLMs, we reformulate trajectory planning as a principle-aligned sequence generation task, decoupling behavior learning from planning principle alignment. Specifically, we first pretrain an autoregressive model to capture the multimodal distribution of expert demonstrations, and then fine-tune the ego policy using reinforcement learning guided by interpretable, rule-based rewards. Experiments on the nuPlan benchmark demonstrate that Plan-R1 achieves SOTA performance, especially in the reactive simulation. Our results underscore the effectiveness of aligning planning trajectory with safety and feasibility objectives through reinforcement learning.